OpenStudy (anonymous):

please help find the general form of the equation of the line that passes through the given point and has the indicated slope. (-2,0), m=-2/5

OpenStudy (jhannybean):

With \((-2~,~0)~,~ m=-\dfrac{2}{5}\) Use point slope form: \(y -y_1 = m(x-x_1)\)

OpenStudy (jhannybean):

So \((x_1~,~y_1) = (-2~,~0)\). Plug that into point slope form.

OpenStudy (jhannybean):

Then tell me what you get :)

OpenStudy (anonymous):

(y-0)= 2/5 (x+2)

OpenStudy (anonymous):

now what?

OpenStudy (jhannybean):

So the general form of the equation follows the format : \(ax +by +c=0\) Right now, we have : \(y=\dfrac{2}{5}(x+2)\) So I guess first we should distribute the \(\dfrac{2}{5}\). What do you have then?

OpenStudy (anonymous):

y= 2/5x+4/5

OpenStudy (jhannybean):

Awesome, now just move everything to the left side.

OpenStudy (jhannybean):

Subtract \(-\dfrac{2}{5}x\) and \(-\dfrac{4}{5}\) from both sides of the equation. What do you end up with?

OpenStudy (anonymous):

2/5x+4/5+y???

OpenStudy (anonymous):

is that right, wrong................?

OpenStudy (jhannybean):

Not quite. you had \(y=\dfrac{2}{5}x+\dfrac{4}{5}\) and you want to fit the general form of the equation: \(ax+by+c=0\), so We would move everything over to the left side by subtracting \(-\dfrac{2}{5}x-\dfrac{4}{5}\) from both sides of he equation. We would end up with: \[\boxed{y-\dfrac{2}{5}x-\dfrac{4}{5} = 0}\]

OpenStudy (jhannybean):

We can also simplify this by multiplying both sides of the equation by \(5\) to get rid of the fractions.: \[5\left(y-\frac{2}{5}x -\frac{4}{5}=0\right) = \boxed{-2x+5y-4=0}\]

OpenStudy (jhannybean):

Do you understand @Romwil ?

OpenStudy (anonymous):

oh sweet thanks

OpenStudy (anonymous):

my answer sheet said it's all positive, but it's prob wrong